how do i find the number of roots in this equation? 3(2x+6)-5=9x+4
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

how do i find the number of roots in this equation? 3(2x+6)-5=9x+4
-----
Solve it:
6x+18-5 = 9x+4
6x+13 = 9x+4
3x = 9
x = 3
----
Looks like the equation has one root.
Cheers,
Stan H.

Answer by richwmiller(17219)

what is the highest power? largest exponent?